(x^(3/2))+(2/the square root of 3)^2
Answer by image_lif(4)

x^3/2+(2/sqrt of 3)^2=0 your eq'n.
(x^3)^1/2+(4/3)=0 simplify
[x((x)^1/2)]^2=[-4/3]^2 sqr both side of the eq'n
x^2(X)=16/9
x^3=16/9 raise both side to 1/3 to get rid of cube root
x=(16/9)^1/3 simplify then you will get
x=2(2/9)^1/3